Margaret walks to the store using the following path: 0.740 miles west, 0.300 miles north, 0.240 miles east. Assume north to be along the +y-axis and west to be along the -x-axis

whats the direction of displacement vector


1
Expert's answer
2021-07-19T09:48:15-0400

Given:

"{\vec d_1}=-0.740\hat i"

"{\vec d_2}=0.300\hat j"

"{\vec d_3}=0.240\hat i"

The displacement

"{\vec d}={\vec d_1}+{\vec d_2}+{\vec d_3}"

So

"{\vec d}=-0.740\hat i+0.300\hat j+0.240\hat i\\
=-0.500\hat i+0.300\hat j"

The direction of displacement is

"\theta=\tan^{-1}\frac{d_y}{d_x}=\tan^{-1}\frac{0.300}{-0.500}=149^{\circ}"

from positive "x" direction or "31^{\circ}\: \rm N\: of\: W".